OpenStudy (anonymous):

-w+4(w+3)=-12 solve for w

OpenStudy (anonymous):

-w +4w + 12 = -12 3w = -12 -12 w = -24/3 w = -8

OpenStudy (anonymous):

w=-8

OpenStudy (aroub):

-w+4w+12=-12 3w=-12-12 w=-8

OpenStudy (anonymous):

thank youuuu:]]]